OpenStudy (strawberry17):

How do you find the interest? You made a purchase with your credit card for $679. You can pay off up to $305 each month. If the APR on this card is 28.23%, how much total interest will you pay on the purchase? Round the solution to the nearest cent.

OpenStudy (tkhunny):

Any way you slice it, it's only a couple of payments. Just do the leg work. Charge: 679.00 Month 1: 679.00*(1 + 0.2823/12) - 305 = Balance Month 2: You tell me!

OpenStudy (strawberry17):

Are you sure that you have to add a one to the APR/12?

OpenStudy (strawberry17):

because then the balance for month is a negative number..

OpenStudy (strawberry17):

month 1

OpenStudy (tkhunny):

The balance is not negative. The "1 +" is required. It represents the accumulation of the debt, not just the interest. Please demonstrate how you can manage a negative value at Month #1. 1 + 0.2823/12 = 1.023525

OpenStudy (strawberry17):

okay, I guess it is not negative... sorry

OpenStudy (tkhunny):

Your verbal description does not answer the question. 679.00*(1 + 0.2823/12) - 305 = Balance 679.00*(1.023525) - 305 = Balance 694.97 - 305 = Balance 389.97 = Balance I don't see a negative balance.

OpenStudy (strawberry17):

you are right..

OpenStudy (tkhunny):

Okay, now do the second month: Month 2: 389.97*(1 + 0.2823/12) - 305 = Balance2

OpenStudy (strawberry17):

month 2 = 94.14

OpenStudy (tkhunny):

Perfect. One more time and we'll have the answer. Don't pay the whole $305.

OpenStudy (strawberry17):

is it $96.35?

OpenStudy (strawberry17):

94.14 X (.2823/12) = 96.35

OpenStudy (tkhunny):

Right on! Original Cost: 679.00 Total Paid: 305.00 + 305.00 + 96.35 = 706.35 How much interest was paid?

OpenStudy (strawberry17):

$27.35 , okay that was the right answer. Thanks so much for explaining it to me!

OpenStudy (tkhunny):

And you wanted to calculate the interest directly! Not always the best choice. Sometime it is. Good work.